# Banana Trees for the Persistence in Time Series Experimentally

### v1

@@ 4.2 Quasi-periodic Data
[…]
@@
For large $\sigma$, the banana trees of the time series show the features we have seen for unbiased random walks, while for small $\sigma$, they locally behave like biased random walks. For example, for $\sigma=1$, we observe about $37\%$ of the items to be critical, while for $\sigma=0.5$ and $\sigma=0.02$ only about $21\%$ and $1\%$ of the items are critical, on average. Like the number of critical items, also the maximum nesting depth for quasi-periodic signals is similar to that of random walks: it scales like $\log n$ for large $\sigma$ and decreases as $\sigma$ goes to 00. In constrast, the average nesting depth is independent of $n$, ranging from $2$ to $4.7$ with mean $2.4$. The behavior of the trail length is illustrated in Figure 5. As shown in the left panel, the average trail length is about $0.5$, with consistently shorter in-trails than mid-trails, on average. This difference is more pronounced for small $\sigma$, when the quasi-periodic signal behaves more like a biased rather than unbiased random walk. The middle and right panels show how the fraction of internal nodes on the longest trail increases as $\sigma$ goes to 00.
@@ 4.3 Real World Examples
We preprocess the data by adding uniformly distributed noise to ensure all values are distinct, making sure the noise is small enough so it breaks ties but does not otherwise alter the ordering,
Table 3: Structural properties of the banana trees constructed for time series from three databases. The maximum nesting depth is measured relative to the depth we observe for unbiased random walks with the same fraction of critical items.
@@
We observe that the average max nesting depth for the three data sets is only a small factor larger than for unbiased random walks; see the middle three columns in Table 3. The largest maximum nesting depth for PhysioNet data is a bit higher, by a factor $6.3$, but $99\%$ of the time series from this data have maximum nesting depth at most $1.68$ times that for unbiased random walks. The average trail lengths are again similar to those of unbiased random walks—see the last two columns, with standard deviation about $0.05$ throughout—and so are the fractions of items in the longest trails, whose geometric means are $0.8\%,0.008\%,0.1\%$ for the in-trails and $0.8\%,0.008\%,0.2\%$ for the mid-trails in the three data sets, compared with about $0.3\%$ and $0.4\%$ in unbiased random walks.
@@ 5 Discussion
Acknowledgments and Disclosure of Funding
The first author is funded by the Vienna Graduate School on Computational Optimization, FWF project no. W1260-N35. The third author is supported by the Wittgenstein Prize, Austrian Science Fund (FWF), grant no. Z 342-N31, and the DFG Collaborative Research Center TRR 109, Austrian Science Fund (FWF), grant no. I 02979-N35.
@@ Appendix A Maintaining Banana Trees
$o$$p$$q$$r$$s$$t$$u$$v$$w$$x$ $o$$p$$q$$r$$s$$t$$u$$v$$w$$x$
Figure 6: 6(a) Increasing the value of a non-critical item $q$ above that of a neighboring non-critical item $p$ results in a new window spanned by the minimum $p$ and the maximum $q$. 6(b) To update the banana tree, we start at the maximum $o$ closest to $q$ such that $o<p<q$ and, beginning with $\texttt{mid}(o)=f$ and following $\texttt{in}(\cdot)$-pointers (dotted red arrows), we search the last node with value greater than $f(q)$; here we land at the node $u$. Then a new banana $(p,q)$ is inserted, in this case as $\texttt{in}(u)$.
@@
Figure 7: 7(a) An interchange of maxima where the value of $j$ increases above that of $q$. Red trails are in-trails and black trails are mid-trails unless labeled otherwise. In the right case, the in-trail of $B(i,j)$ becomes the mid-trail of $B(i,q)$, and the mid-trail of $B(i,j)$ becomes the in-trail of $B(i,q)$. There are two additional cases that are similar to the reverse of the cases shown here. 7(b) An interchange of minima where the value of $i$ decreases below the value of $p$. The nodes $a$, $b$, and $j$ satisfy $f(a)<f(j)<f(b)$. For nodes in sections of trail between $j$ and $q$ the low pointer is updated to point to $i$.
$a$$b$$j$$k$$a$$b$$j$$k$       $a$$b$$j$$k$$j$$b$$a$$k$
Figure 8: 8(a) Processing an injury for cutting. The green dashed section of the in-trail of $B(j,k)$ is inserted into the mid-trail above $a$. 8(b) Processing a fatality for cutting. The green dashed section of the banana $B(j,k)$ is swapped with $a$. This section consists of the in-trail of $B(j,k)$, the node $j$ and a section of the mid-trail of $B(j,k)$. After the swap the in-trail between $a$ and $k$ is empty, shown as a dotted line.
Banana trees can be constructed in linear time by traversing the critical items from left to right. During the traversal, each item is processed once to be inserted into the banana tree and once to correctly assign low and dth pointers. We refer to [9] for details and focus on the maintenance algorithms.
We first discuss how the banana trees are updated when the value of a single item changes. An arbitrary value change is processed by a sequence of local maintenance operations, which we describe here.
If $j$ is non-critical in $g$, it may become critical in $h(j)$. We distinguish two cases: case (1) where $j$ becomes critical by passing the value of a critical neighbor, i.e., by its value either increasing above a neighboring maximum or decreasing below a neighboring minimum, and case (2) where $j$ becomes critical by passing the value of a non-critical neighbor. In case (1) the banana tree is updated by a slide: the critical neighbor becomes non-critical, and thus its node in the banana tree is simply updated to refer to $j$. In case (2) the item $j$ and its neighbor $k$ both become critical and $j$ either becomes a maximum or a minimum, depending on whether its value increased or decreased. Assume $j$ becomes a maximum. The banana tree is updated by inserting a new banana $(k,j)$. To find where to insert the new banana the algorithm begins at the maximum $m$ such that $m<k<j$ or $j<k<m$. If $j<m$ it follows the leftmost path beginning at the child of $m$ immediately to its right; if $j>m$ it follows the rightmost path beginning at the child of $m$ immediately to its left. The node $j$ is inserted below the last node $d$ such that $f(d)>f(j)$. See the dotted path in Figure 6.
If $j$ is critical in $g$, it may become non-critical in $h(j)$. It may either (a) become non-critical by passing the value of a non-critical neighbor—which is the reverse of case (1) above—or (b) become non-critical by passing the value of a critical neighbor—which is the reverse of case (2) above. In case (a) the banana tree is also updated by a slide (see above); in case (b) the critical neighbor of $j$ also becomes non-critical. They will span a banana when this happens and we need to only delete it to update the banana tree. We call this operation a cancellation.
Finally, an item $j$ that is critical may remain critical in $h$, but changing its value may affect the order of critical items. We first consider the case where $j$ is a maximum. The banana tree only needs to be updated if the value of $j$ increases above $\texttt{up}(j)$, or if it decreases below $\texttt{mid}(j)$, $\texttt{in}(j)$ or $\texttt{dn}(j)$ (whichever has greatest value). In this case we perform an interchange of maxima, which resembles a rotation and which is illustrated in Figure 7(a). Note that this operation only reassigns a constant number of pointers. The second case to consider is the one where $j$ is a minimum. The banana tree needs to be updated if either the value of $j$ falls below the value of $\texttt{low}(\texttt{dth}(j))$, or rises above the value of a node $k$ with $j=\texttt{low}(\texttt{dth}(k))$. This is achieved by an interchange of minima, which is illustrated in Figure 7(b).
@@
We first describe how to cut a banana tree between items $j$ and $j+1$ at a virtual item $c=j+\frac{1}{2}$. The cases $f(j)>f(j+1)$ and $f(j)<f(j+1)$ are symmetric and we assume the former. First, items $p$ and $q$ are inserted such that $j<p<q<j+1$ and $f(j)>f(q)>f(p)>f(j+1)$, setting $f(q)=f(p)+\varepsilon$ with $\varepsilon>0$ such that $p$ and $q$ are paired. The items $p$ and $q$ become the new endpoints of the left and right interval, respectively, and are inserted into the banana trees via an anti-cancellation. The bananas where $c$ is one of the three panels of the corresponding window are stored in a stack, with $B(p,q)$ on the bottom and value of the maximum increasing towards the top. Each banana $B(a,b)$ on the stack is classified according to the position of $c$ in relation to $a$ and $b$: if $c<a<b$ or $b<a<c$, then $c$ lies in the in-panel, and we say that $B(a,b)$ experiences an injury; if $a<c<b$ or $b<c<a$, then $c$ lies in the mid-panel, and we say that $B(a,b)$ experiences a fatality; if $a<b<c$ or $c<b<a$, then $c$ lies in the out-panel, and we say that $B(a,b)$ experiences a scare. To perform actual cutting an empty tree is constructed consisting of a special root and a dummy node. The bananas on the stack are then processed from top to bottom. Based on injury, fatality or scare nodes are moved between the two trees, until $B(p,q)$ has been processed and the two trees correspond to the left and right list. Figure 8 illustrates the processing of injuries and fatalities. Scares are processed via an interchange of minima.
@@
Appendix B Implementation Details
[…]
Figure 9: 9(a) A window spanned by minimum $u$ and maximum $p$ on the left and the corresponding section of the banana tree. Rectangles indicate windows with the in-panel and mid-panel separated by a dotted line. 9(b) An internal node $a$ and a leaf node $b$ that span a banana $B(b,a)$ and pointers defined at these nodes. The pointer $\texttt{low}(b)=b$ is not shown.
Our implementation of banana trees is based on pointers between nodes, as illustrated in Figure 9(b). Each node has pointers in and mid, which point to one end of the doubly-linked lists that represent the in-trail and mid-trail. The nodes on a trail are connected by pointers up and dn, which point to the next node with higher and lower value, respectively. Each minimum stores a pointer dth, which points to the maximum that it is paired with. Note that if $\texttt{dth}{b}=a$, then nodes $b$ and $a$ are the ends of an in-trail and a mid-trail. This establishes the pairing in the persistence diagram. Finally, every node stores a pointer low that points to the lower end of the trail containing the node, or the node itself if it is a minimum. In the example in Figure 9(a), $\texttt{low}{r}=\texttt{low}{t}=b$. The pointer low can be used to compute the inverse of the pointer dth. For example, given $\texttt{dth}(d)=c$ we have $\texttt{low}(\texttt{in}(c))=\texttt{low}(\texttt{mid}(c))=d$. Furthermore, every node stores a pointer to the item it represents.
We implemented the banana trees in C++20, using boost.intrusive [15] for the binary search trees. Specifically, we use splay trees [22] whenever we require splitting or joining, and AVL trees [2] otherwise. We exploit the fact that the down-tree is just the up-tree for $-f$ and implement banana trees as a template class, where the sole parameter is the sign used for accessing $-f$. Key to the performance of splitting and joining is the use of a central memory pool for allocating nodes and items. This moves ownership of nodes and items out of the classes representing banana trees and intervals, which eliminates the need to move items between individual instances when splitting or joining.
Appendix C Additional Experiments
C.1 Random Walks
@@
Figure 10: Timing the construction of banana trees. 10 Slowdown of banana trees compared to Gudhi for unbiased random walks ($\mu=0$). For each number of items, $n$, the ribbon extends from the minimum to the maximum observed slowdown, and the point shows the geometric mean. 10 Average number of allocations by the memory pools for storing banana tree nodes and items. 10 Slowdown compared to Gudhi for different values of the parameter $\mu$ in the generation of random walks. The ribbon extends from the minimum to the maximum observed slowdown, and the point shows the geometric mean.
@@
Figure 11: 11 The maximum nesting depth of the leaf bananas in the up-tree of an unbiased random walk ($\mu=0$) of size $n$. The ribbon extends from the minimum to the maximum obsered average nesting depth, and the dot shows the mean. The red line is the graph of a function of the form $\beta_{1}\log n$ fit via linear regression. 11 Maximum nesting depth of 100 up-trees for biased random walks of size $2\cdot 10^{6}$. The ribbon extends from the minimum to the maximum per bias parameter, $\mu$, and the dot shows the mean.
C.2 Linear-Time Cases
C.2.1 Definitions
@@
The linear-time case for topological maintenance is a function $w_{t}:\mathbb{Z}\to\mathbb{R}$ defined over $n$ items for some odd $n>4$, which we give below. First, we define an auxiliary function
@@
$w_{t}(i)$ $=-s(\left\lfloor\frac{n}{2}\right\rfloor-1-i)$ for $\qquad\text{for }0\leq i<\left\lfloor\frac{n}{2}\right\rfloor,$ (5)
$w_{t}(i)$ $=s(i-\left\lfloor\frac{n}{2}\right\rfloor)+0.1$ for $\qquad\text{for }\left\lfloor\frac{n}{2}\right\rfloor\leq i<n.$

### v2

@@ 4.2 Quasi-Periodic Data
$r(0)=0\mbox{\rm\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }r(i)=r(i-1)+\eta_{i}\mbox{\rm\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ }1\leq i<n,$ (2)
@@
For large $\sigma$, the banana trees of the time series show the features we have seen for unbiased random walks, while for small $\sigma$, they locally behave like biased random walks. For example, for $\sigma=1$, we observe about $37\%$ of the items to be critical, while for $\sigma=0.5$ and $\sigma=0.02$ only about $21\%$ and $1\%$ of the items are critical, on average. Like the number of critical items, also the maximum nesting depth for quasi-periodic signals is similar to that of random walks: it scales like $\log n$ for large $\sigma$ and decreases as $\sigma$ goes to 00. In contrast, the average nesting depth is independent of $n$, ranging from $2$ to $4.7$ with mean $2.4$. The behavior of the trail length is illustrated in Figure 6. As shown in the left panel, the average trail length is about $0.5$, with consistently shorter in-trails than mid-trails, on average. This difference is more pronounced for small $\sigma$, when the quasi-periodic signal behaves more like a biased rather than unbiased random walk. The middle and right panels show how the fraction of internal nodes on the longest trail increases as $\sigma$ goes to 00.
@@ 4.3 Real World Examples
We preprocess the data by adding uniformly distributed noise to ensure all values are distinct, making sure the noise is small enough so it breaks ties but does not otherwise alter the ordering.
Table 3: Structural properties of the banana trees constructed for time series from three databases. The average, minimum, and maximum of the maximum nesting depth is measured relative to the depth we observe for unbiased random walks with the same fraction of critical items.
@@
We observe that the average max nesting depth for the three data sets is only a small factor larger than for unbiased random walks; see the middle three columns in Table 3. The largest maximum nesting depth for PhysioNet data is a bit higher, by a factor $6.3$, but $99\%$ of the time series from this data have maximum nesting depth at most $1.68$ times that for unbiased random walks. The average trail lengths are again similar to those of unbiased random walks—see the last two columns, with standard deviation about $0.05$ throughout—and so are the fractions of items in the longest trails, whose geometric means are $0.8\%,0.008\%,0.1\%$ of all critical items for the in-trails and $0.8\%,0.008\%,0.2\%$ of all critical items for the mid-trails in the three data sets, compared with about $0.3\%$ and $0.4\%$ in unbiased random walks.
@@ Appendix A Maintaining Banana Trees
[…]
Figure 7: 7(a) A window spanned by minimum $u$ and maximum $p$ on the left and the corresponding section of the banana tree. Rectangles indicate windows with the in-panel and mid-panel separated by a dotted line. 7(b) An internal node $a$ and a leaf node $b$ that span a banana $B(b,a)$ and pointers defined at these nodes. The pointer $\texttt{low}(b)=b$ is not shown.
Banana trees are described by pointers between nodes, as illustrated in Figure 7(b). Each node has pointers in and mid, which point to one end of the doubly-linked lists that represent the in-trail and mid-trail. The nodes on a trail are connected by pointers up and dn, which point to the next node with higher and lower value, respectively. Each minimum stores a pointer dth, which points to the maximum that it is paired with. Note that if $\texttt{dth}{b}=a$, then nodes $b$ and $a$ are the ends of an in-trail and a mid-trail. This establishes the pairing in the persistence diagram. Finally, every node stores a pointer low that points to the lower end of the trail containing the node, or the node itself if it is a minimum. In the example in Figure 7(a), $\texttt{low}(r)=\texttt{low}(t)=b$. The pointer low can be used to compute the inverse of the pointer dth. For example, given $\texttt{dth}(d)=c$ we have $\texttt{low}(\texttt{in}(c))=\texttt{low}(\texttt{mid}(c))=d$. Furthermore, every node stores a pointer to the item it represents.
$m$$k$$j$$r$$s$$t$$u$$v$$w$$x$ $m$$k$$j$$r$$s$$t$$u$$v$$w$$x$
Figure 8: 8(a) Increasing the value of a non-critical item $j$ above that of a neighboring non-critical item $k$ results in a new window spanned by the minimum $k$ and the maximum $j$. 8(b) To update the banana tree, we start at the maximum $m$ closest to $j$ such that $m<k<j$ and, beginning with $\texttt{mid}(m)=w$ and following $\texttt{in}(\cdot)$-pointers (dotted red arrows), we search the last node with value greater than $f(j)$; here we land at the node $u$. Then a new banana $B(k,j)$ is inserted, in this case with $\texttt{in}(u)$ being set to point to $j$.
Banana trees can be constructed in linear time by traversing the critical items from left to right. During the traversal, each item is processed once to be inserted into the banana tree and once to assign pointers representing the nesting of bananas and the pairing of items. We refer to [8] for details and focus on the maintenance algorithms.
We first discuss how the banana trees are updated when the value of a single item changes. An arbitrary value change is processed by a sequence of local maintenance operations, which we describe here. Let $f:[1,m]\to\mathbb{R}$ and $g:[1,m]\to\mathbb{R}$ be the maps before and after the update, respectively. We assume that the value of item $j$ is changed and $f(j)\neq g(j)$. The maps $f$ and $g$ agree in the values of all other items, i.e., $f(i)=g(i)$ for all $i\neq j$.
If $j$ is non-critical in $f$, it may become critical in $g$. We distinguish two cases: case (1) where $j$ becomes critical by passing the value of a critical neighbor, i.e., by its value either increasing above a neighboring maximum or decreasing below a neighboring minimum, and case (2) where $j$ becomes critical by passing the value of a non-critical neighbor. In case (1) the banana tree is updated by a slide: the critical neighbor becomes non-critical, and thus its node in the banana tree is simply updated to refer to $j$. In case (2) the item $j$ and its neighbor $k$ both become critical and $j$ either becomes a maximum or a minimum, depending on whether its value increased or decreased. Assume $j$ becomes a maximum as illustrated in Figure 8. The banana tree is updated by inserting a new banana $B(k,j)$. To find where to insert the new banana the algorithm begins at the maximum $m$ such that $m<k<j$ or $j<k<m$. If $j<m$ it follows the leftmost path beginning at the child of $m$ immediately to its right; if $j>m$ it follows the rightmost path beginning at the child of $m$ immediately to its left. The node $j$ is inserted below the last node $d$ such that $f(d)>f(j)$. See the dotted path in Figure 8.
If $j$ is critical in $f$, it may become non-critical in $g$. It may either (a) become non-critical by passing the value of a non-critical neighbor—which is the reverse of case (1) above—or (b) become non-critical by passing the value of a critical neighbor—which is the reverse of case (2) above. In case (a) the banana tree is also updated by a slide (see above); in case (b) the critical neighbor of $j$ also becomes non-critical. They will span a banana when this happens and we only need to only this banana to update the banana tree. This operation called a cancellation.
@@
Figure 9: 9(a) An interchange of maxima where the value of $j$ increases above that of $q$. Red trails are in-trails and black trails are mid-trails unless labeled otherwise. In the right case, the in-trail of $B(i,j)$ becomes the mid-trail of $B(i,q)$, and the mid-trail of $B(i,j)$ becomes the in-trail of $B(i,q)$. There are two additional cases that are similar to the reverse of the cases shown here. 9(b) An interchange of minima where the value of $i$ decreases below the value of $p$. The nodes $a$, $b$, and $j$ satisfy $f(a)<f(j)<f(b)$. For nodes in sections of trail between $j$ and $q$ the low pointer is updated to point to $i$.
Finally, an item $j$ that is critical may remain critical in $g$, but changing its value may affect the order of critical items. We first consider the case where $j$ is a maximum. The banana tree only needs to be updated if the value of $j$ increases above $\texttt{up}(j)$, or if it decreases below $\texttt{mid}(j)$, $\texttt{in}(j)$ or $\texttt{dn}(j)$ (whichever has greatest value). In this case we perform an interchange of maxima, which resembles a rotation and which is illustrated in Figure 9(a). Note that this operation only reassigns a constant number of pointers. The second case to consider is the one where $j$ is a minimum. The banana tree needs to be updated if either the value of $j$ falls below the value of $\texttt{low}(\texttt{dth}(j))$, or rises above the value of a node $k$ with $j=\texttt{low}(\texttt{dth}(k))$. This is achieved by an interchange of minima, which is illustrated in Figure 9(b).
@@
$a$$b$$j$$k$$a$$b$$j$$k$       $a$$b$$j$$k$$j$$b$$a$$k$
Figure 10: 10(a) Processing an injury for cutting. The green dashed section of the in-trail of $B(j,k)$ is inserted into the mid-trail above $a$. 10(b) Processing a fatality for cutting. The green dashed section of the banana $B(j,k)$ is swapped with $a$. This section consists of the in-trail of $B(j,k)$, the node $j$ and a section of the mid-trail of $B(j,k)$. After the swap the in-trail between $a$ and $k$ is empty, shown as a dotted line.
@@
We first describe how to cut a banana tree between items $j$ and $j+1$ at a virtual item $c=j+\frac{1}{2}$. The cases $f(j)>f(j+1)$ and $f(j)<f(j+1)$ are symmetric and we assume the former. First, items $p$ and $q$ are inserted such that $j<p<q<j+1$ and $f(j)>f(q)>f(p)>f(j+1)$, setting $f(q)=f(p)+\varepsilon$ with $\varepsilon>0$ such that $p$ and $q$ are paired. The items $p$ and $q$ become the new endpoints of the left and right interval, respectively, and are inserted into the banana trees via an anti-cancellation. The bananas where $c$ lies in one of the three panels of the corresponding window are stored in a stack, with $B(p,q)$ on the bottom and the value of the maximum increasing towards the top of the stack. Each banana $B(a,b)$ on the stack is classified according to the position of $c$ in relation to $a$ and $b$: if $c<a<b$ or $b<a<c$, then $c$ lies in the in-panel, and we say that $B(a,b)$ experiences an injury; if $a<c<b$ or $b<c<a$, then $c$ lies in the mid-panel, and we say that $B(a,b)$ experiences a fatality; if $a<b<c$ or $c<b<a$, then $c$ lies in the out-panel, and we say that $B(a,b)$ experiences a scare. To perform actual cutting an empty tree is constructed consisting of a special root and a dummy node. The bananas on the stack are then processed from top to bottom. Based on injury, fatality or scare nodes are moved between the two trees, until $B(p,q)$ has been processed and the two trees correspond to the left and right list. Figure 10 illustrates the processing of injuries and fatalities. Scares are processed via an interchange of minima.
@@
Appendix B Additional Experiments
B.1 Random Walks
@@
Figure 11: Timing the construction of banana trees. 11 Slowdown of banana trees compared to Gudhi for unbiased random walks ($\mu=0$). For each number of items, $n$, the ribbon extends from the minimum to the maximum observed slowdown, and the point shows the geometric mean. 11 Average number of allocations by the memory pools for storing banana tree nodes and items. 11 Slowdown compared to Gudhi for different values of the parameter $\mu$ in the generation of random walks. The ribbon extends from the minimum to the maximum observed slowdown, and the point shows the geometric mean.
@@
Figure 12: 12 The maximum nesting depth of the leaf bananas in the up-tree of an unbiased random walk ($\mu=0$) of size $n$. The ribbon extends from the minimum to the maximum obsered average nesting depth, and the dot shows the mean. The red line is the graph of a function of the form $\beta_{1}\log n$ fit via linear regression. 12 Maximum nesting depth of 100 up-trees for biased random walks of size $2\cdot 10^{6}$. The ribbon extends from the minimum to the maximum per bias parameter, $\mu$, and the dot shows the mean.
B.2 Linear-Time Cases
B.2.1 Definitions
@@
The linear-time case for topological maintenance is a function $w_{t}:\mathbb{Z}\to\mathbb{R}$ defined over $n$ items, with $n$ a multiple of $4$, which we give below. First, we define an auxiliary function
@@
$w_{t}(i)$ $=-s\left(\frac{n}{2}-1-i\right)$ for $\qquad\text{for }0\leq i<\frac{n}{2},$ (5)
$w_{t}(i)$ $=s\left(i-\frac{n}{2}\right)-0.1$ for $\qquad\text{for }\frac{n}{2}\leq i<n.$